OpenStudy (anonymous):

In△ABC the m∠A=(2x+23)?, m∠B=(3x−10)?,and m∠C=(2x+34)?. Choose the correct statement about the classification of the triangle by the angle measures. A. The triangle is a right triangle because x = 19 , and there is one right angle in the triangle. B. The triangle is an obtuse triangle because x = 19 , and there is one angle that is greater than 90 degrees. C. The triangle is an acute triangle because x = 19 , and all three angles are less than 90 degrees. D. The triangle is an equiangular triangle because x = 19 , and all the angles are congruent to each other.

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

Over here, you see, you actually have to put in the values given in each option

OpenStudy (anonymous):

OK.

OpenStudy (perl):

can you upload a diagram

OpenStudy (anonymous):

It doesn't include one.

OpenStudy (anonymous):

Yeah. That should do it? the options have also defined what are the different kinds of triangles. should you face any more problems? :-)

OpenStudy (anonymous):

OK, I'll let you know.

OpenStudy (anonymous):

You wont need a diagram right @perl . Cause the angles are given. We gotta check with values of x in the options and write the answer accordingly

OpenStudy (anonymous):

It's not A

OpenStudy (anonymous):

Yes correct :-) Go on..

OpenStudy (anonymous):

It's C?

OpenStudy (perl):

we know that a triangle adds up to 180 degrees

OpenStudy (perl):

<A + < B + C = (2x+23)+ (3x−10)+ (2x+34) = 180

OpenStudy (perl):

<A + < B + C = (2x+23)+ (3x+10)+ (2x+34) = 180

OpenStudy (anonymous):

Right

OpenStudy (perl):

we can solve it by algebra

OpenStudy (perl):

collect like terms

OpenStudy (anonymous):

I'll solve it.

OpenStudy (anonymous):

I got 113/7 o_o

OpenStudy (anonymous):

is it right ?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

what happens when you plug in x = 19

OpenStudy (perl):

m∠A=(2x+23), m∠B=(3x−10)?,and m∠C=(2x+34)

OpenStudy (perl):

m<A=(2x+23) m<A = 2*19 + 23 m∠B=3*19−10 m∠C=2*19+34

OpenStudy (anonymous):

SO IT'S c?